OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 4y = –9 2x + 5y = –6

OpenStudy (anonymous):

using substitution x + 4y = –9 x= –9-4y substitute it to the second equation

OpenStudy (anonymous):

4x + 5(–4x – 2) = –18 4x + 5(4x – 2) = –18 4(–4x – 2) + 5y = –18 4(4x – 2) + 5y = –18

OpenStudy (anonymous):

wich 1

OpenStudy (anonymous):

which answer is it???

OpenStudy (anonymous):

out of the four answerrs i gave you wich 1